OpenStudy (aye_my_peeps):

evaluate:

OpenStudy (aye_my_peeps):

\[\frac{ 3*1.5-2.5 }{ 4(8.62-6.37) }\]

OpenStudy (aveline):

What do you need help with? On the top, you use order of operations. On the bottom, you use the distributive property.

OpenStudy (aye_my_peeps):

bottom

OpenStudy (aveline):

Ok, 4(8.62-6.37)=(4*8.62)-(4*6.37)

OpenStudy (welshfella):

or , on the bottom you can work out the operation in the parentheses first then multiply by 4.

OpenStudy (welshfella):

thats a bit easier

OpenStudy (aye_my_peeps):

\[\frac{ 2 }{ 9 }\]

OpenStudy (aveline):

Correct ^^

OpenStudy (aye_my_peeps):

thnx :)